# Epistemic Uncertainty in Conformal Scores: A Unified Approach

### v1

@@ B.3 Base model hyperparameters
Appendix C EPICSCORE Computational Details
This section outlines the specifications of each predictive model used within our framework.
C.1 Splitting strategy
When splitting the calibration set $\mathcal{D}_{\text{cal}}$ for fitting the predictive model and deriving EPICSCORE’s adaptive cutoffs in disjunct data subsets, we prioritize allocating the majority of data to model training, as cutoff derivation primarily involves a simpler quantile computation. Specifically, for small with $n\leq 3000$, we reserve 30% of the calibration samples for cutoff computation. For larger datasets, this allocation is capped at 1000 samples to maintain computational efficiency. This approach ensures that EPICSCORE achieves both accurate predictive distribution estimates and well-calibrated cutoffs.
C.2 MDN MC-Dropout details
@@
Optimization relies on the Adam optimizer [Kingma, 2014] with a learning rate of 0.001 and a step learning rate scheduler that decays by 0.99 every 5 epochs. To model epistemic uncertainty, we incorporate MC dropout [Gal and Ghahramani, 2016], applying a dropout rate of 0.5 to each hidden layer. All weights are initialized using PyTorch’s default settings. To monitor generalization performance, we reserve 30% of the training data exclusively for validation. The model is trained for up to 2,000 epochs, with early stopping triggered after 50 epochs of no improvement on the validation set to prevent overfitting. Additionally, we apply feature scaling and target normalization to enhance numerical stability and improve parameter estimation.
@@
C.3 Variational GP details
For the Variational Gaussian Process (GP), we implement the model using the gpytorch package [Gardner et al., 2018] and PyTorch. We first define the GP prior with a constant mean function and an RBF kernel for the covariance. To approximate the posterior, we employ a Natural Variational Distribution [Salimbeni et al., 2018], using 15 inducing points for small datasets ($n<10000$) and $50$ for medium and large-scale data. Training is performed by minimizing the negative variational ELBO, where the Natural Gradient Descent [Salimbeni et al., 2018] updates the variational parameters, while the Adam optimizer [Kingma, 2014] refines the GP kernel and noise variance hyperparameters.
@@ C.3 Variational GP details
C.4 BART details
For the Bayesian Additive Regression Tree (BART) model, we use the base implementation from pymc3 [Quiroga et al., 2022] and adopt the heteroscedastic variant [Pratola et al., 2020]. The conformal scores are modeled as a normal distribution, where the mean is determined by the sum of regression trees, and the variance depends on $\mathbf{X}$. We set the number of trees to 100, while keeping the prior hyperparameters at their default values in pymc3. Target normalization is applied to enhance numerical stability and improve posterior estimation.
@@ C.4 BART details
Appendix D Proofs
@@
By a similar argument, we also have:
@@
Therefore, under the event $\Omega$ of probability at least $1-\delta$, we conclude that

### v2

@@ C.3 Base model hyperparameters
Appendix D EPICSCORE Computational and Methodological Details
This section outlines the specifications of each predictive model used in our framework, discusses the scalability of each approach, and details the trade-offs associated with prior choices.
D.1 Splitting strategy
When splitting the calibration set $\mathcal{D}_{\text{cal}}$ for fitting the predictive model and deriving EPICSCORE’s adaptive cutoffs in disjunct data subsets, we prioritize allocating the majority of data to model training, as cutoff derivation primarily involves a simpler quantile computation. Specifically, for small with $n\leq 3000$, we use 30% of the calibration samples for cutoff computation. For larger datasets, this allocation is capped at 1000 samples to maintain computational efficiency. This approach ensures that EPICSCORE achieves both accurate predictive distribution estimates and well-calibrated cutoffs.
D.2 MDN MC-Dropout details
@@
Optimization relies on the Adam optimizer [Kingma, 2014] with a learning rate of 0.001 and a step learning rate scheduler that decays by 0.99 every 5 epochs. To model epistemic uncertainty, we incorporate MC dropout [Gal and Ghahramani, 2016], applying a dropout rate of 0.5 to each hidden layer. This setup implicitly corresponds to a diffuse, non-informative prior over the network weights—one that does not favor any specific parameter configuration and assumes equal uncertainty over weights. All network weights are initialized using PyTorch’s default scheme. To monitor generalization, 30% of the training data is held out for validation. The model is trained for up to 2,000 epochs, with early stopping after 50 epochs without validation improvement. Feature scaling and target normalization are also applied to improve numerical stability and parameter estimation.
@@
D.3 Variational GP details
For the Variational Gaussian Process (GP), we implement the model using the gpytorch package [Gardner et al., 2018] and PyTorch. The GP prior is specified with a constant mean function and a Radial Basis Function (RBF) kernel, which provides a smoothness assumption but remains essentially non-informative in the sense that it does not encode strong prior beliefs about the function values or parameters. To approximate the posterior, we employ a Natural Variational Distribution [Salimbeni et al., 2018], using 15 inducing points for small datasets ($n<10000$) and $50$ for medium and large-scale data. Training is performed by minimizing the negative variational ELBO, where the Natural Gradient Descent [Salimbeni et al., 2018] updates the variational parameters, while the Adam optimizer [Kingma, 2014] refines the GP kernel and noise variance hyperparameters.
@@ D.3 Variational GP details
D.4 BART details
For the Bayesian Additive Regression Tree (BART) model, we use the base implementation from pymc3 [Quiroga et al., 2022] and adopt the heteroscedastic variant [Pratola et al., 2020]. The conformal scores are modeled as a normal distribution, where the mean is determined by the sum of regression trees, and the variance depends on $\mathbf{X}$. We configure the model with 100 trees and retain the default prior settings provided by pymc3, which correspond to a non-informative prior favoring deeper trees. This choice allows for flexible modeling without imposing strong structural constraints. Additionally, we apply target normalization to improve numerical stability and enhance posterior inference.
@@ D.4 BART details
D.5 Impact of Prior Choices
In this section, we illustrate the impact of prior choices in EPICSCORE through a simulated experiment comparing the prediction intervals produced by EPICSCORE-BART under two prior configurations: diffuse and concentrated. To emulate data sparsity, we follow a similar setup to Section A.1, simulating a scenario with both high- and low-density regions. Specifically, given a sample size $n$, we draw $\left\lfloor{0.8\cdot n}\right\rfloor$ samples from $X\sim U(-1,0)$, representing the denser region. The remaining $\left\lfloor{0.2\cdot n}\right\rfloor$ samples are generated from $X\sim\text{Beta}(5,5)$, concentrating points in the intermediate region of the interval [0,1]. The response variable follows a heteroscedastic distribution $Y\sim N\left(\mathbb{I}\{x>0\},0.05+(\sin^{2}(15\cdot x)\cdot\mathbb{I}\{x>0\})\right)$, which results in low aleatoric and epistemic uncertainty over the densely sampled interval $[-1,0]$ and high uncertainty—both epistemic and aleatoric—over the sparse interval $[0,1]$.
The dataset is divided into training, calibration, and test sets, each containing 200 samples. As the base predictor, we use a k-nearest neighbors (k-NN) regression model with $k=10$, trained on the training set. For EPICSCORE, both BART configurations share the same settings: $m=50$ trees, heteroscedastic variance, a normal likelihood, response normalization, and $1000$ Monte Carlo samples. The key difference lies in the prior hyperparameter $\beta$,which governs tree depth by controlling the probability that a node is non-terminal. A higher value, such as $\beta=0.9$ yields a diffuse prior with deeper trees and greater model flexibility, while a lower value like $\beta=0.1$ results in a concentrated prior with shallower trees, imposing stronger regularization.
Figure 7 illustrates how the prediction intervals behave under each prior configuration.
Figure 7: Comparison of prediction regions produced by EPICSCORE using BART under two prior configurations: a diffuse prior and a concentrated prior. While the diffuse prior produces wider prediction regions—particularly in sparse areas—due to greater model flexibility, the concentrated prior yields narrower intervals reflecting stronger regularization. This highlights the sensitivity of EPICSCORE to prior specification.
As shown, the diffuse prior yields wider intervals—particularly in data-sparse regions—capturing heightened epistemic uncertainty. In contrast, the concentrated prior produces tighter intervals in denser regions, reflecting its stronger regularization and more confident assumptions in well-populated areas. This highlights the sensitivity of EPICSCORE to prior specification and demonstrates how prior concentration directly influences the adaptiveness and sharpness of the prediction sets.
To further investigate this trade-off, we also evaluate the influence of $\beta$ using the AISL metric, which jointly accounts for coverage and interval length. As shown in Figure 8, higher values of $\beta$ lead to deeper trees and wider intervals, while smaller values yield more concise intervals due to stronger regularization.
Figure 8: Variation of the Average Interval Score Loss (AISL) as a function of the $\beta$ hyperparameter in the BART model used within the EPICSCORE framework. Lower AISL values indicate more favorable balance, suggesting that tuning $\beta$ can significantly affect performance by controlling model flexibility and prior regularization.
This analysis helps reveal the trade-off between interval regularization and accurate representation of epistemic uncertainty. Crucially, it also provides a foundation for tuning prior hyperparameters more systematically. By optimizing AISL—or another relevant performance metric—one can select prior configurations in a principled manner. In more complex scenarios involving multiple hyperparameters, this process remains scalable through the use of Bayesian optimization techniques [Frazier, 2018].
D.6 Scalability comparisson
Incorporating Bayesian models into conformal prediction offers a principled way to capture epistemic uncertainty but may introduce additional computational overhead. To quantify this cost within our framework, we evaluate the runtime of EPICSCORE across increasing sample sizes using the three Bayesian predictive models discussed previously: MDN, BART, and Variational GP. We consider a synthetic regression task with 30 input features and one response variable, generated using scikit-learn’s make_regression function [Pedregosa et al., 2011]. For each sample size ($n=1000,2000,5000,10000,20000,50000$), we generate training and calibration sets of equal size ($n$), train a k-nearest neighbors (k-NN) model with $k=30$ as the base predictor on the training data, and fit the predictive models on the calibration data. Each setting is repeated 10 times, and Figure 9 reports the average runtime along with its standard deviation for each sample size.
Figure 9: Computational cost of EPICSCORE as a function of sample size $n$ (log scale), using the three Bayesian predictive models: BART, MDN and Variational GP. While all three models show comparable running times at moderate scales(e.g., $n=10000$), MDNs exhibit significantly better scalability and faster execution as $n$ increases, underscoring their advantage in large-scale settings.
The MDN configuration includes two hidden layers (64–64 units), ReLU activations, 50% dropout, $2000$ training epochs, a learning rate of $0.001$, and batch size dynamically scaled to the sample size (ranging from $35$ to $450$). Inputs are standardized, and the response is normalized. BART is configured with $50$ trees, normal heteroscedastic errors, $500$ posterior samples, and parallelization using $6$ CPU cores. The GP model is trained via a variational approximation, with the number of inducing points dynamically adjusted from $15$ to $150$ based on $n$, a variational learning rate of $0.1$, a hyperparameter learning rate of $0.01$, and identical batch size and early stopping parameters to the MDN.
Overall, MDNs demonstrate the most stable and scalable runtime across all sample sizes. Variational Gaussian Processes also scale well, though they are generally less efficient than MDNs. The observed drop in runtime between $5000$ and $10000$ samples for both neural network–based models likely stems from the early stopping mechanism, which halts training once convergence is detected. BART, on the other hand, maintains low computational cost for small datasets but exhibits a sharp increase in runtime as the sample size grows. This highlights a key trade-off: while BART is effective for small to medium-scale problems, neural network–based models like MDNs are better suited for larger datasets, where the iterative nature of MCMC-based methods such as BART becomes increasingly expensive.
This experiment was conducted using an AMD Ryzen 7 6000 series CPU. For the largest dataset, which includes over $100000$ calibration observations, the runtime for each method remained small: BART completed in under 16 minutes, MDN in under 2 minutes, and GP in under 4 minutes. These results highlight the computational efficiency of the proposed approaches, particularly the scalability of MDNs and GPs for large-scale applications.
D.7 Application to density-based conformity scores
In this section, we use a simulated experiment to illustrate how EPICSCORE can enhance density-based conformal prediction methods by explicitly incorporating epistemic uncertainty. Following the same simulation setup described in Section D.5, we compare our approach to a standard method for conformal prediction with a density-based score, namely HPD-split [Izbicki et al., 2022]. The experiment is designed to reflect regions with both high epistemic and aleatoric uncertainty, allowing us to highlight the advantages of our method in capturing complex uncertainty structures.
The base conditional density model is a Mixture Density Network (MDN) consisting of two hidden layers with 64 and 128 neurons, respectively, and a Gaussian mixture output with 5 components. The model incorporates a dropout rate of 0.15 to improve generalization and is trained on 2000 samples. For EPICSCORE, we use a Bayesian Additive Regression Trees (BART) model with 50 trees, a gamma error distribution, heteroscedastic variance, and a non-informative prior. The tree depth is controlled by a fixed hyperparameter $\beta=0.9$, promoting deeper trees. Calibration is performed using 1000 posterior samples. The HPD-split prediction region is constructed directly from the MDN’s conditional density estimates. Figure 10 compares the prediction regions obtained by the two methods.
Figure 10: Comparison of prediction regions generated by the HPD (Highest Predictive Density) split [Izbicki et al., 2022] and EPICSCORE methods in a simulated scenario. Both methods produce similar regions in the high-density area ($x<0$), but EPICSCORE exhibits sharper adaptivity in the sparse region ($x>0$), offering more informative and expanded predictive sets where epistemic uncertainty is higher.
Compared to the HPD-based region, our method exhibits significantly better adaptiveness and uncertainty quantification, particularly in data-sparse regions of the input space. At the same time, it preserves tight prediction intervals in denser, data-rich areas. This demonstrates EPICSCORE’s ability to more effectively capture uncertainty in challenging settings where data is limited, without sacrificing performance in well-populated regions.
Appendix E Proofs
@@
By a symmetric argument, we also have:
@@
Therefore, conditionally to the event $\Omega$, of probability at least $1-\delta$, we conclude that